OpenStudy (anonymous):

wats a^2+b^2 formula????

OpenStudy (anonymous):

\[(a + b)^{2} = a ^{2} +b ^{2} + 2ab\] \[Therefore, a ^{2} +b ^{2} = a ^{2} +b ^{2} + 2ab - 2ab = (a + b)^{2} - 2ab\]

OpenStudy (vishweshshrimali5):

\[(a+b)^2=a^2+b^2+2ab\] \[(a+b)^2-2ab=a^2+b^2\] \[(a-b)^2=a^2+b^2-2ab\] \[(a-b)^2+2ab=a^2+b^2\]